A cannon shoots a bullet from the roof of a building from the position (in SI units): π‘ŸΜ‚0 = 9 𝑖̂+ 4 𝑗̂+ 15 π‘˜Μ‚ and with initial velocity (SI units): 𝑣̂o = 13 𝑖̂+ 22.5 𝑗̂+ 15 π‘˜Μ‚ where the z-axis points upward in the vertical direction and with origin at the ground level. Assuming that the air resistance is negligible, find the maximum height reached by the bullet and the position where it hits the ground.


1
Expert's answer
2021-09-20T10:00:57-0400
"h=15+15t-0.5(9.8)t^2\\H=26.5\ m\\t_0=3.855\ s\\
x=9+13(3.855)=59.1\ m\\y=4+22.5(3.855)=90.7\ m\\"